OpenStudy (anonymous):

A bag of fertilizer covers 2500 square feet of lawn. Find how many bags of fertilizer should be purchased to cover a rectangular lawn 410 feet by 65 feet.

OpenStudy (jhannybean):



OpenStudy (jhannybean):

\[\ A_R = x\cdot y\] or \[\ A_R= l\cdot w\] where l = length and w= width.

OpenStudy (jhannybean):

They give you \(\ A_R = 2500 \ ft^2\)

OpenStudy (jhannybean):

\( x= 410 \ ft\) and \( y = 65 \ ft\) \(\ \therefore\) \[\ A_R = x\cdot y = (410 \ ft)(65 \ ft)\]